OpenStudy (anonymous):

Giuliana has 22 quarters and dollar coins worth a total of $10.75. Which answer works? 0.1n + 0.05(n – 175) = 13.30 0.1n + 0.05(175 – n) = 13.30 0.1(n – 175) + 0.05 = 13.30 0.1(175 – n) + 0.05n = 13.30

OpenStudy (anonymous):

HEEELP

OpenStudy (ybarrap):

q=# quarters d=# dollars n=# coins n=q+d \(\implies\) q=n-d q*.25+d*1=10.75 \(\implies\)(n-d)*.25+d=10.75 \(\text{ after substituting}\) = (22-d)*.25+d=10.75 From which you can find the number of dollars, \(d\)/ Number of quarters then is q=22-d Not of the equations in your list matches this problem. Does this make sense?

OpenStudy (anonymous):

So it's the second one right?